Good Day to you all. Could you please solve the following questions for me:
1). 20m 4dm divided by 12
2). 12cm 16mm divided by 8
3). 9km 300m divided by 4
: Good Day to you all. Could you please solve the following questions for me:
1). 20m 4dm divided by 12
2). 12cm 16mm divided by 8
3). 9km 300m divided by 4

Answer by Fombitz(1789)

The key is to convert to equivalent units then divide.
.
.
.
1. 20m 4dm divided by 12
20m=2m(100 dm/m)=200dm
20m 4dm=204 dm
204 dm/12=17 dm
.
.
.
2.12cm 16mm=12(10)+16=136mm
136mm/8=17 mm
.
.
.
3.9 km 300m=9(1000)+300=9300 m
9300m/4=2325 m